# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ Appendix B Detailed Pro-LNN Architecture and Training Procedure
Step 1. Stable liquid cell. At the core of Pro-LNN is a stable liquid cell that updates a hidden state $\mathbf{h}_{i}^{(\nu)}\in\mathbb{R}^{d}$ given the current input $\mathbf{z}_{i}^{(\nu)}$ and the previous state. Concretely, we first compute neuron-wise time constants and a pre-activation: $\bm{\tau}^{(\nu)}=\mathrm{softplus}\big(W_{\tau}[\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu)}]\big)+\tau_{\min}$ and $\tilde{\mathbf{u}}^{(\nu)}=W_{h}\mathbf{h}_{i}^{(\nu)}+W_{x}\mathbf{z}_{i}^{(\nu)}+\mathbf{b}$, where $W_{\tau},W_{h},W_{x}$ and $\mathbf{b}$ are trainable, and $\tau_{\min}>0$ enforces strictly positive time scales (we also clamp $\bm{\tau}^{(\nu)}$ to $\tau_{\min},\tau_{\max}]$ for numerical stability). The effective update rate is then $\bm{\alpha}^{(\nu)}=\mathrm{clip}\Big(\frac{\Delta\nu}{\bm{\tau}^{(\nu)}},\alpha_{\min},\alpha_{\max}\Big)$, with a fixed $\Delta\nu$ (one trading interval in our implementation) and small constants $0<\alpha_{\min}\leq\alpha_{\max}<1$. We then apply a lightweight layer normalization and non-linearity, $\tilde{\mathbf{h}}_{i}^{(\nu)}=\tanh\!\big(\mathrm{LN}(\tilde{\mathbf{u}}^{(\nu)})\big)$, and update the hidden state as
@@
Notation. For clarity, all nonlinear operators in the above update are applied element-wise unless otherwise stated. Specifically, $\mathrm{softplus}(\cdot)$ denotes the standard softplus activation $\mathrm{softplus}(x)=\log(1+e^{x})$, $\mathrm{clip}(x,a,b)$ clips each element of $x$ into the interval $[a,b]$, $\mathrm{LN}(\cdot)$ denotes layer normalization [24], and $\mathrm{LiquidCell}(\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu)})$ is a shorthand for one step of the stable liquid update in Step 1 (using the time–constant gating and residual integration described above).
Step 2. Historical encoding. Given the latest $L$ observations of ES $s_{i}$ up to trading round $\nu$, Pro-LNN encodes the history by iteratively applying the liquid cell: $\mathbf{h}_{i}^{(\nu-L+1)}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(\nu-L+1)},\mathbf{0}\big),\ldots,\mathbf{h}_{i}^{(\nu)}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu-1)}\big)$.
@@
Step 3. Multi-horizon readout and consistency. After encoding the history, a shallow MLP readout maps the final hidden state to an $H$-step forecast:
@@ 21  Update parameters $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$;
Step 4. Inference and integration. At inference, Pro-LNN takes, for each ES $s_{i}$, the latest $L$ observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, encodes them via the liquid cell into $\mathbf{h}_{i}^{(\nu)}$, and applies the readout to obtain an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. These predictions are computed in the offline stage before each trading round and then fed into: (i) the eACO-VRP module for time-window-aware ES recommendation and AgP route pre-planning, and (ii) the auction module for ES–AgP contract design. Since Pro-LNN runs offline, its wall-clock inference latency does not affect the real-time trading between SDs and SPs, while its compact footprint and robustness to noisy inputs make it well suited for deployment on edge and near-edge computing platforms.
Appendix C Proof of Property of Off-AIC2
Theorem 2.
@@ Appendix C Proof of Property of Off-AIC2
All winning ESs and AgPs obtained from Off-AIC2 are individually rational during the offline stage.
@@
The individual rationality of Off-AIC2 is guaranteed by the core design of its allocation (Phase 2) and pricing (Phase 3) procedures in Alg. 2, under the single-parameter assumption that each ES $s_{i}$ has a per-unit valuation and each AgP $v_{k}$ has a per-unit cost.
@@
$\bullet$ Allocation in Phase 2: For each AgP $v_{k}$, the candidate ES set is constructed as
@@
where $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$ denotes the surplus of matching ES $s_{i}$ with AgP $v_{k}$. Hence, if ES $s_{i}$ is eventually matched (i.e., there exists some $k$ such that $x^{\mathrm{Off}}_{i,k}=1$), the corresponding pair $(i,k)$ must satisfy
@@
Under truthful bidding, $\mathrm{bid}_{i,k}$ equals $s_{i}$’s true per-unit valuation for the contracted service. Therefore, already at the allocation stage, every winning ES enjoys a strictly positive surplus margin between its valuation and the AgP’s ask.
$\bullet$ Pricing in Phase 3: In Phase 3, for each winning ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$, Off-AIC2 treats its effective bid as a variable and performs a monotone binary search over an interval of the form
@@
where $\overline{\mathrm{bid}}_{i}$ is $s_{i}$’s reference bid (coinciding with its true valuation when truthful) and $\underline{b}_{i}$ is a market-wide lower bound. The binary search stops at the minimal critical payment that keeps $s_{i}$ in the winning set when Phase 2 is re-executed. By construction, we have
@@
(ii) AgP side (sellers). We now turn to AgP sellers $v_{k}$.
$\bullet$ Allocation in Phase 2: In Phase 2, AgPs only serve ESs that yield strictly positive surplus, and they are protected by capacity constraints. Specifically, $v_{k}$ may serve ES $s_{i}$ only if $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}>0$, and the matching obeys
$\sum_{i}x^{\mathrm{Off}}_{i,k}N_{i}^{\mathrm{Dem}}\leq C_{k}^{\mathrm{V}}.$ (32)
Under truthful reporting, $\mathrm{ask}_{k}$ equals $v_{k}$’s true per-unit cost. Thus each winning AgP only accepts contracts where the buyer-side valuation exceeds its cost, ensuring a nonnegative surplus margin at the allocation stage.
$\bullet$ Pricing in Phase 3: In Phase 3, for each winning AgP $v_{k}$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$, Off-AIC2 performs a monotone binary search over an interval
@@
where $\mathrm{ask}_{k}$ is its own ask (true cost when truthful) and $\overline{r}_{k}$ is a market-determined upper bound. The search stops at the maximal critical reward that still keeps $v_{k}$ in the winning set when Phase 2 is rerun. Consequently,
@@
$u_{k}^{\mathrm{AgP}}=\big(r_{k}^{\mathrm{s}*}-\mathrm{ask}_{k}\big)\cdot\Big(\sum_{i}x^{\mathrm{Off}}_{i,k}N_{i}^{\mathrm{Dem}}\Big)\;\geq\;0.$ (35)
In summary, by construction of the candidate sets and critical-value pricing, every winning ES pays no more than its valuation, and every winning AgP receives no less than its cost. Therefore, Off-AIC2 is individually rational for both ESs and AgPs in the offline stage. ∎
Theorem 3.
Under the single-parameter assumption, the proposed Off-AIC2 mechanism is truthful (dominant-strategy incentive-compatible) for both ESs and AgPs in the offline stage.
@@
The truthfulness of Off-AIC2 follows from the fact that, on both the ES side and the AgP side, it implements a monotone allocation rule together with critical-value payments, which is sufficient for dominant-strategy incentive compatibility in single-parameter environments.
(i) ES side (buyers). Fix any ES $s_{i}$ and treat all other bids and asks as given. Under the single-parameter assumption, $s_{i}$ submits a single per-unit bid $b_{i}$, which induces its bid entries $\mathrm{bid}_{i,k}=b_{i}$ and average bid $\overline{\mathrm{bid}}_{i}=b_{i}$.
$\bullet$ Monotone allocation (Phases 1 & 2). In Phase 1, ESs are sorted in nonincreasing order of $\overline{\mathrm{bid}}_{i}=b_{i}$. Increasing $b_{i}$ can only move $s_{i}$ upward in this list; it can never push $s_{i}$ to a worse rank. Since Off-AIC2 only considers the top-$I^{\ast}$ ESs, once $s_{i}$ belongs to the top-$I^{\ast}$ at some bid $b_{i}$, it remains in the top-$I^{\ast}$ for any higher bid $b^{\prime}_{i}>b_{i}$.
In Phase 2, a necessary condition for $s_{i}$ to be matched is that it lies in the top-$I^{\ast}$ ESs and that there exists an AgP $v_{k}$ such that
$U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}=b_{i}-\mathrm{ask}_{k}>0,\qquad\mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}.$ (36)
If $s_{i}$ is selected at some bid $b_{i}$, then increasing its bid to $b^{\prime}_{i}>b_{i}$ strictly increases $U_{i,k}$ for all $k$ and does not violate any of the above inequalities. Under a fixed tie-breaking rule in the eACO-VRP procedure, any ES that is winning at $b_{i}$ remains winning at any higher bid $b^{\prime}_{i}>b_{i}$. Thus the ES-side allocation rule is monotone in $b_{i}$.
$\bullet$ Critical payment (Phase 3). In Phase 3, for each winning ES $s_{i}$ (with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$), Off-AIC2 treats its effective bid as a variable and performs a monotone binary search over an interval
@@
repeatedly re-running Phase 2 to check whether $s_{i}$ would still be selected. The search stops at the smallest bid at which $s_{i}$ remains a winner, and the resulting value is taken as its payment:
$p_{i}^{\mathrm{b}*}=\text{critical bid of }s_{i}.$ (38)
By construction, this critical payment depends only on other participants’ bids and asks and on the system parameters; it does not depend on the actual report $b_{i}$ as long as $b_{i}\geq p_{i}^{\mathrm{b}*}$. Consequently, $s_{i}$ cannot reduce its payment by overbidding, and any underbid $b_{i}<p_{i}^{\mathrm{b}*}$ either keeps the allocation and payment unchanged (if $s_{i}$ still wins) or makes it lose a profitable contract. Hence truthful bidding maximizes $s_{i}$’s utility, and Off-AIC2 is truthful for ESs.
@@
(ii) AgP side (sellers). We now fix an AgP $v_{k}$ and consider its per-unit ask $a_{k}$, treating all ES bids and other AgPs’ asks as fixed. Under the single-parameter assumption, $a_{k}$ is $v_{k}$’s single strategic parameter.
@@
$\bullet$ Monotone allocation (Phases 1 & 2). In Phase 1, AgPs are sorted in nondecreasing order of $\mathrm{ask}_{k}=a_{k}$. Lowering $a_{k}$ can only move $v_{k}$ upward in this list, never downward. Since Off-AIC2 only considers the top-$K^{\ast}$ AgPs, once $v_{k}$ is in the top-$K^{\ast}$ at some ask $a_{k}$, it remains in the top-$K^{\ast}$ for any lower ask $a^{\prime}_{k}<a_{k}$.
In Phase 2, $v_{k}$’s attractiveness to ESs is determined by the surplus terms
$U_{i,k}=\mathrm{bid}_{i,k}-a_{k}.$ (39)
If $v_{k}$ is selected at some ask $a_{k}$, then lowering its ask to $a^{\prime}_{k}<a_{k}$ strictly increases all $U_{i,k}$, making $v_{k}$ (weakly) more competitive. Under the same capacity constraint and tie-breaking rule in eACO-VRP, this can never cause $v_{k}$ to lose its winning status. Thus, once $v_{k}$ is a winner at $a_{k}$, it remains a winner at any lower ask $a^{\prime}_{k}<a_{k}$, so the seller-side allocation rule is monotone in $a_{k}$.

### v2

@@ B-B Detailed Pro-LNN Architecture and Training Procedure
Step 1. Stable liquid cell (line 7, Alg. 1): At the core of Pro-LNN is a stable liquid cell that updates a hidden state $\mathbf{h}_{i}^{(\nu)}\in\mathbb{R}^{d}$ given the current input $\mathbf{z}_{i}^{(\nu)}$ and the previous state. Concretely, we first compute neuron-wise time constants and a pre-activation: $\bm{\tau}^{(\nu)}=\mathrm{softplus}\big(W_{\tau}[\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu)}]\big)+\tau_{\min}$ and $\tilde{\mathbf{u}}^{(\nu)}=W_{h}\mathbf{h}_{i}^{(\nu)}+W_{x}\mathbf{z}_{i}^{(\nu)}+\mathbf{b}$, where $W_{\tau},W_{h},W_{x}$ and $\mathbf{b}$ are trainable, and $\tau_{\min}>0$ enforces strictly positive time scales (we also clamp $\bm{\tau}^{(\nu)}$ to $\tau_{\min},\tau_{\max}]$ for numerical stability). The effective update rate is then $\bm{\alpha}^{(\nu)}=\mathrm{clip}\Big(\frac{\Delta\nu}{\bm{\tau}^{(\nu)}},\alpha_{\min},\alpha_{\max}\Big)$, with a fixed $\Delta\nu$ (one trading interval in our implementation) and small constants $0<\alpha_{\min}\leq\alpha_{\max}<1$. We then apply a lightweight layer normalization and non-linearity, $\tilde{\mathbf{h}}_{i}^{(\nu)}=\tanh\!\big(\mathrm{LN}(\tilde{\mathbf{u}}^{(\nu)})\big)$, and update the hidden state as
@@
Notation. For clarity, all nonlinear operators in the above update are applied element-wise unless otherwise stated. Specifically, $\mathrm{softplus}(\cdot)$ denotes the standard softplus activation $\mathrm{softplus}(x)=\log(1+e^{x})$, $\mathrm{clip}(x,a,b)$ clips each element of $x$ into the interval $[a,b]$, $\mathrm{LN}(\cdot)$ denotes layer normalization [24], and $\mathrm{LiquidCell}(\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu)})$ is a shorthand for one step of the stable liquid update in Step 1 (using the time-constant gating and residual integration described above).
Step 2. Historical encoding (lines 4-7, Alg. 1): Given the latest $L$ observations of ES $s_{i}$ up to trading round $\nu$, Pro-LNN encodes the history by iteratively applying the liquid cell: $\mathbf{h}_{i}^{(\nu-L+1)}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(\nu-L+1)},\mathbf{0}\big),\ldots,\mathbf{h}_{i}^{(\nu)}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu-1)}\big)$.
@@
Step 3. Multi-horizon readout and consistency (lines 8-17, Alg. 1): After encoding the history, a shallow MLP readout maps the final hidden state to an $H$-step forecast:
@@ 21  Update parameters $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$;
Step 4. Inference and integration (line 9, Alg. 1): At inference, Pro-LNN takes, for each ES $s_{i}$, the latest $L$ observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, encodes them via the liquid cell into $\mathbf{h}_{i}^{(\nu)}$, and applies the readout to obtain an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. These predictions are computed in the offline stage before each trading round and then fed into: (i) the eACO-VRP module for time-window-aware ES recommendation and AP route pre-planning, and (ii) the auction module for ES–AP contract design. Since Pro-LNN runs offline, its wall-clock inference latency does not affect the real-time trading between SDs and SPs, while its compact footprint and robustness to noisy inputs make it well suited for deployment on edge and near-edge computing platforms.
B-C Detailed pseudo-code of eACO-VRP
In this section, we provide the detailed pseudo-code of eACO-VRP in Alg. 2.
[…]
2
[…]
4
5while $iter\leq iter_{\mathrm{max}}$ do
6  for each ant $\mathbbm{k}\in\mathbb{K}$ do
7     $\bm{M}^{\mathbbm{k}}\!\leftarrow\!\varnothing$;  $\mathrm{n}_{r}\!\leftarrow\!\mathrm{n}_{0}$
8     while end condition not met do
9        Compute: feasible set $J^{\mathbbm{k}}(\mathrm{n}_{r})$ and transition probabilities $pr^{\mathbbm{k}}_{r,s}$ (via (6))
10        if $J^{\mathbbm{k}}(\mathrm{n}_{r})\neq\varnothing$ then
11           sample next node $\mathrm{n}_{s}$ according to $p^{\mathbbm{k}}_{r,s}$ from $J^{\mathbbm{k}}(\mathrm{n}_{r})$
12           $\bm{M}^{\mathbbm{k}}\!\leftarrow\!\bm{M}^{\mathbbm{k}}\cup\{\text{edge}(\mathrm{n}_{r},\mathrm{n}_{s})\}$
13           $\mathrm{n}_{r}\!\leftarrow\!\mathrm{n}_{s}$
14       else
15          break
16
17
18
19 $\bm{M}^{\mathrm{max}}\!\leftarrow\!\arg\max_{\bm{M}^{\mathbbm{k}}:\mathbbm{k}\in\mathbb{K}}\;\sum_{(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{W}}_{r,s}}$
[…]
21    Pheromone evaporation and deposit (7) and (8)
22
23 $iter\!\leftarrow\!iter+1$
24
Algorithm 2 Proposed eACO-VRP
B-D Detailed pseudo-code of Off-AIC2
In this section, we provide the detailed pseudo-code of Off-AIC2 in Alg. 3.
1Input: ES set $\bm{\mathcal{S}}$ with demands $\{N_{i}^{\mathrm{Dem}}\}$ and per-ES bids $\{\overline{\mathrm{bid}}_{i}\}$; AP set $\bm{\mathcal{V}}$ with asks $\{\mathrm{ask}_{k}\}$ and capacities $\{C_{k}^{\mathrm{A}}\}$
2Initialization: Set $\bm{X}^{*}\!\leftarrow\!\bm{0}$, $\bm{p}^{\mathrm{b}*}\!\leftarrow\!\bm{0}$, $\bm{r}^{\mathrm{s}*}\!\leftarrow\!\bm{0}$.
3Sort ES indices into $\mathcal{L}^{\mathrm{ES}}$ by $\overline{\mathrm{bid}}_{i}$ (descending), and AP indices into $\mathcal{L}^{\mathrm{AP}}$ by $\mathrm{ask}_{k}$ (ascending); relabel $(i,k)$ accordingly.
4 Find a pair of pivotal ranks $(I^{\ast},K^{\ast})$ such that $\overline{\mathrm{bid}}_{I^{\ast}}\geq\mathrm{ask}_{K^{\ast}},\quad\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}\;\geq\;\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$, and $I^{\ast}$ is maximized (breaking ties by larger capacity slack $\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}-\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$).
5Initialize the set of unmatched ESs: $\mathcal{U}^{\mathrm{free}}\leftarrow\{s_{i}\mid 1\leq i\leq I^{\ast}\}$.
6 for $k$ in top-$K^{\ast}$ APs (according to $\mathcal{L}^{\mathrm{AP}}$) do
7  if $C_{k}^{\mathrm{A}}\leq 0$ or $\mathcal{U}^{\mathrm{free}}=\emptyset$ then
8     continue
9 Build candidate ES set $\mathcal{C}_{k}\leftarrow\big\{s_{i}\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}^{\mathrm{Off}}>0\ \wedge\ \overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}\big\}$;
10 if $\mathcal{C}_{k}=\emptyset$ then
11     continue
12 $\mathcal{K}_{k}^{\star}\leftarrow$ Alg. 2 ;
13  for $i\in\mathcal{K}_{k}^{\star}$ do
14     Set $x^{\mathrm{Off}}_{i,k}\leftarrow 1$;
15     Remove $s_{i}$ from $\mathcal{U}^{\mathrm{free}}$: $\mathcal{U}^{\mathrm{free}}\leftarrow\mathcal{U}^{\mathrm{free}}\setminus\{s_{i}\}$;
16 Update remaining capacity: $C_{k}^{\mathrm{A}}\leftarrow C_{k}^{\mathrm{A}}-\sum_{i\in\mathcal{K}_{k}^{\star}}N_{i}^{\mathrm{Dem}}$.
17Denote the resulting matching by $\bm{X}^{*}=\{x^{\mathrm{Off}}_{i,k}\}$.
18for each matched ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$ do
19  Set $high\leftarrow\overline{\mathrm{bid}}_{i}$, $low\leftarrow\overline{\mathrm{bid}}_{I^{\ast}+1}$;
20  while $low<high$ do
21     Set $b_{\mathrm{eff}}\leftarrow(low+high)/2$;
22     Rerun Step 2 on the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs, using $b_{\mathrm{eff}}$ as ES $s_{i}$’s bid and keeping all other bids/asks unchanged, to obtain a temporary matching $\bm{X}^{\mathrm{Temp}}$;
23     if $\sum_{k}x^{\mathrm{Temp}}_{i,k}=1$ then $high\leftarrow b_{\mathrm{eff}}$;
24     else $low\leftarrow b_{\mathrm{eff}}$;
25
26 Set $p_{i}^{\mathrm{b}*}\leftarrow high$.
27
28for each matched AP $v_{k}$ with $\sum_{s_{i}}x^{\mathrm{Off}}_{i,k}\geq 1$ do
30  while $low<high$ do
31     Set $\mathrm{ask}_{\mathrm{eff}}\leftarrow(low+high)/2$;
32     Rerun Step 2 on the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs, using $\mathrm{ask}_{\mathrm{eff}}$ as AP $k$’s ask (and keeping all other bids/asks unchanged), to obtain a temporary matching $\bm{X}^{\mathrm{Temp}}$;
33     if $\sum_{i}x^{\mathrm{Temp}}_{i,k}\geq 1$ then $low\leftarrow\mathrm{ask}_{\mathrm{eff}}$;
34     else $high\leftarrow\mathrm{ask}_{\mathrm{eff}}$;
35
36 Set $r_{k}^{\mathrm{s}*}\leftarrow low$.
37
Output: Matching $\bm{X}^{*}$, ES payments $\bm{p}^{\mathrm{b}*}$, AP rewards $\bm{r}^{\mathrm{s}*}$.
Algorithm 3 Proposed Off-AIC2
B-E Economic Properties of Off-AIC2
Proposition 1.
(Individual Rationality) An auction is said to be individually rational if the payment made by any winning ES (to its matched AP) does not exceed its bid, and the revenue obtained by any winning AP is no less than its asked price.
Proposition 2.
(Near-Truthfulness) An auction mechanism is said to be near-truthful if every ES and AP maximizes, or almost maximizes, its utility by reporting its true private valuation. In particular, for any fixed bids and asks of the other participants, no single ES or AP can increase its utility by more than a small margin through unilaterally misreporting its bid or ask. In this sense, truthful reporting remains an empirically near-optimal strategy, and large deviations from true valuations are not beneficial.
Proposition 3.
(Budget Balance) An auction satisfies budget balance if the total payment collected from all winning ESs is no less than the total payment made to all winning APs.
Proposition 4.
@@ B-E Economic Properties of Off-AIC2
All winning ESs and APs obtained from Off-AIC2 are individually rational during the offline stage.
The proposed offline stage of FUSION is shown to meet all the above properties, with detailed theoretical analysis as follow.
@@
The individual rationality of Off-AIC2 is guaranteed by the core design of its allocation (Step 2) and pricing (Step 3) procedures in Alg. 3, under the single-parameter assumption that each ES $s_{i}$ has a per-unit valuation and each AP $v_{k}$ has a per-unit cost.
@@
$\bullet$ Allocation in Step 2: For each AP $v_{k}$, the candidate ES set is constructed as
@@
where $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$ denotes the surplus of matching ES $s_{i}$ with AP $v_{k}$. Hence, if ES $s_{i}$ is eventually matched (i.e., there exists some $k$ such that $x^{\mathrm{Off}}_{i,k}=1$), the corresponding pair $(s_{i},v_{k})$ must satisfy
@@
Under truthful bidding, $\mathrm{bid}_{i,k}$ equals $s_{i}$’s true per-unit valuation for the contracted service. Therefore, already at the allocation stage, every winning ES enjoys a strictly positive surplus margin between its valuation and the AP’s ask.
$\bullet$ Pricing in Step 3: In Step 3, for each winning ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$, Off-AIC2 treats its effective bid as a variable and performs a monotone binary search over an interval of the form
@@
where $\overline{\mathrm{bid}}_{i}$ is $s_{i}$’s reference bid (coinciding with its true valuation when truthful) and $\underline{b}_{i}$ is a market-wide lower bound. The binary search stops at the minimal critical payment that keeps $s_{i}$ in the winning set when Step 2 is re-executed. By construction, we have
@@
(ii) AP side (sellers). We now turn to AP sellers $v_{k}$.
$\bullet$ Allocation in Step 2: In Step 2, APs only serve ESs that yield strictly positive surplus, and they are protected by capacity constraints. Specifically, $v_{k}$ may serve ES $s_{i}$ only if $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}>0$, and the matching obeys
$\sum_{i}x^{\mathrm{Off}}_{i,k}N_{i}^{\mathrm{Dem}}\leq C_{k}^{\mathrm{A}}.$ (32)
Under truthful reporting, $\mathrm{ask}_{k}$ equals $v_{k}$’s true per-unit cost. Thus each winning AP only accepts contracts where the buyer-side valuation exceeds its cost, ensuring a nonnegative surplus margin at the allocation stage.
$\bullet$ Pricing in Step 3: In Step 3, for each winning AP $v_{k}$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$, Off-AIC2 performs a monotone binary search over an interval
@@
where $\mathrm{ask}_{k}$ is its own ask (true cost when truthful) and $\overline{r}_{k}$ is a market-determined upper bound. The search stops at the maximal critical reward that still keeps $v_{k}$ in the winning set when Step 2 is rerun. Consequently, we have
@@
$u_{k}^{\mathrm{AP}}=\big(r_{k}^{\mathrm{s}*}-\mathrm{ask}_{k}\big)\cdot\Big(\sum_{i}x^{\mathrm{Off}}_{i,k}N_{i}^{\mathrm{Dem}}\Big)\;\geq\;0.$ (35)
In summary, by construction of the candidate sets and critical-value pricing, every winning ES pays no more than its valuation, and every winning AP receives no less than its cost. Therefore, Off-AIC2 is individually rational for both ESs and APs in the offline stage. ∎
Proposition 5 (Near-truthfulness of Off-AIC2).
Off-AIC2 mechanism is near-truthful for both ESs and APs in the offline stage.
@@
At a high level, Off-AIC2 follows the same structural design as classical truthful single-parameter auctions: on both the ES side and the AP side, it combines a (approximately) monotone allocation rule with critical-value-like payments obtained via binary search. In fully monotone settings, such a combination is sufficient for dominant-strategy incentive compatibility; in our case, the presence of route planning via eACO-VRP makes the allocation only approximately monotone, and therefore we obtain approximate (rather than exact) truthfulness.
(i) ES side (buyers). Fix any ES $s_{i}$ and treat all other bids and asks as given. Under the single-parameter simplification, $s_{i}$ reports a single per-unit bid $b_{i}$, which induces its bid row $\mathrm{bid}_{i,k}=b_{i}$ and average bid $\overline{\mathrm{bid}}_{i}=b_{i}$.
Approximate monotone allocation (Steps 1 & 2). In Step 1, ESs are sorted in nonincreasing order of $\overline{\mathrm{bid}}_{i}=b_{i}$. Increasing $b_{i}$ can only move $s_{i}$ upward in this list; it never pushes $s_{i}$ to a worse rank. Since Off-AIC2 only considers the top-$I^{\ast}$ ESs, once $s_{i}$ belongs to the top-$I^{\ast}$ at some bid $b_{i}$, it remains in the top-$I^{\ast}$ for any higher bid $b^{\prime}_{i}>b_{i}$.
In Step 2, a necessary condition for $s_{i}$ to be matched is that it lies in the top-$I^{\ast}$ ESs and that there exists an AP $v_{k}$ such that
$U_{i,k}^{\mathrm{Off}}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}=b_{i}-\mathrm{ask}_{k}>0,\qquad\overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}.$ (36)
If we ignore the internal randomness and heuristic nature of eACO-VRP, raising $b_{i}$ strictly increases all $U_{i,k}^{\mathrm{Off}}$, making $s_{i}$ (weakly) more attractive and therefore easier to be selected. In practice, the eACO-VRP-based matching can occasionally break strict monotonicity because ES–AP matching and route planning are coupled in a combinatorial way; however, for a fixed tie-breaking rule and fixed routes of other APs, the allocation rule behaves approximately monotone in $b_{i}$: once $s_{i}$ is selected at some bid level, it typically remains selected for any sufficiently close higher bid $b^{\prime}_{i}>b_{i}$.
Critical-value-like payment (Step 3). In Step 3, for each winning ES $s_{i}$ (with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$), Off-AIC2 treats its effective bid as a variable and performs a monotone binary search over an interval
@@
repeatedly re-running Step 2 to check whether $s_{i}$ would still be selected. The search stops at the smallest bid at which $s_{i}$ remains a winner, and the resulting value is taken as its payment $p_{i}^{\mathrm{b}*}$. When the allocation subroutine is exactly monotone, $p_{i}^{\mathrm{b}*}$ coincides with the classical critical value and yields exact dominant-strategy truthfulness. With eACO-VRP, the binary-search procedure still emulates such a critical value and produces a payment that depends only weakly on $s_{i}$’s own report as long as $b_{i}$ stays in a small neighborhood above $p_{i}^{\mathrm{b}*}$. Our numerical results (Fig. 2) confirm that $s_{i}$’s realized utility curve is single-peaked around its true valuation, and deviations from truthful bidding only bring marginal utility gains, if any. Hence truthful bidding is a locally optimal strategy and nearly utility-maximizing for ESs.
@@
(ii) AP side (sellers). Now fix an AP $v_{k}$ and consider its per-unit ask $a_{k}$, treating all ES bids and other APs’ asks as fixed. Under the single-parameter simplification, $a_{k}$ is $v_{k}$’s single strategic parameter.
@@
Approximate monotone allocation (Steps 1 & 2). In Step 1, APs are sorted in nondecreasing order of $\mathrm{ask}_{k}=a_{k}$. Lowering $a_{k}$ can only move $v_{k}$ upward in this list, never downward. Since Off-AIC2 only considers the top-$K^{\ast}$ APs, once $v_{k}$ lies in the top-$K^{\ast}$ at some ask $a_{k}$, it remains in the top-$K^{\ast}$ for any lower ask $a^{\prime}_{k}<a_{k}$.
In Step 2, $v_{k}$’s attractiveness to ESs is determined by the surplus terms
$U_{i,k}^{\mathrm{Off}}=\mathrm{bid}_{i,k}-a_{k}.$ (38)
Lowering $a_{k}$ strictly increases all $U_{i,k}^{\mathrm{Off}}$, making $v_{k}$ more competitive. With a fixed eACO-VRP configuration, this behavior again yields an allocation rule that is approximately monotone in $a_{k}$: once $v_{k}$ is selected at some ask $a_{k}$, it tends to remain selected for any sufficiently close lower ask $a^{\prime}_{k}<a_{k}$, except for rare non-monotone cases introduced by heuristic route decisions.